Statement of Income Year Ended 31 December 2018 Total sales Operating costs and expenses EBITDA Depreciation Operating incomeThe target capital structure for Acme is 60% debt and 40% equity. The risk-free rate is 2%. The market risk premium is 5%. The beta of Acme against a relevant stock index is 0.8. The pre-tax cost of debt for Acme is 4%. Assume an effective corporation tax of 17%.

Question 1
(a) Based on the given financial statements, calculate the following ratios for 2018. Explain the consequence if a calculated ratio is significantly lower than its corresponding industry average.
(i) Quick ratio
(ii) Average collection period
(iii) Fixed asset turnover
(iv) Operating profit margin
(v) Times interest earned ratio
(vi) Degree of financial leverage (DFL)
(b) Compute Acme Corporation’s return on equity for 2018 using the Du Pont technique.

The following is a set of financial statements for Acme Corporation Financial Statements for Acme Corporation (in thousands, except for Per-Share Data) Year Ended 31-Dec Balance Sheet Assets Current assets Cash andequivalents Accounts receivable Inventory Statement of Cash Flows Year Ended 31 December Operating activities Net income Adjustments 2017 2018 2018 332 190 560 410 292 Depreciation Changes in working capital Accounts receivable Inventories Accounts payable Accrued taxes and expenses Cash provided by operating activities 300 Total current assets 1,160 1,332 2,200 2,600 1,200 1,3001,400 2,460 2,732 (30) 15 10 587 Gross fixed assets Accumulated depreciation 900 Net fixed assets Total assets Liabilities and shareholders' equity Current liabilities Accounts payable Notes payable Accrued taxes and expenses Investing activities Purchases of fixed assets 300 250 150 Cash used for investing activities 140 625 865 100 200 Financing activities Notes payable Long-term financing issuances Common stock dividends 50 25 160 85 Total current liabilities Long-term debt Common stock Additional paid-in capital Retained earnings 890 100 Cash used for financing activities 842 1,142 2,732 102 190 292 Cash and equivalents increase (decrease) Cash and equivalents at beginning of year Cash and equivalents at end of year Supplemental cash flow disclosures Interest paid Income taxes paid Total shareholders' equity 970 2,460 Total liabilities and shaeholders' equity 100 Statement of Income Year Ended 31 December 2018
Statement of Income Year Ended 31 December 2018 Total sales Operating costs and expenses EBITDA Depreciation Operating income (EBIT) Interest expense Income before tax Taxes (at 17% Net income Dividends Change in retained earnings Earnings per share (EPS) Dividends per share 3,000 2,200 800 300 500 100 400 68 332 160 172 0.664 $0.320
Data for new project under consideration $1,000,000 5 Investment in new machinery Useful life of the machinery (years) Expected life of the project (years) Salein the first year from new machine implementation COGS as a percentage of sales Estimated selling price at the the end of 3 years Additional working capital is 20% of sales attributed to new machine $800,000 70% $450,000 $160,000
sales is expected to increase 15% per year from the new machine project. Working capital at the beginning of a given year is 20% of sales for that year, allocated at the beginning of the project and recouped at the end of project life. Depreciation uses straight line method and reduces the machine salvage value to zero over its useful life. The CFO is confident of this project and expects the company-wide dividends to grow at the same rate as the sales growth of this new project for the next 3 years before reverting to a constant perpetual rate of 49 Acme Corporation equity rights issuance data Current share price Suggested subscription price $20 $15
Answer #1

a) Ratio calculations for 2018:

i) Quick ratio =( Current assets - Inventory )/ Current liabilities

Quick ratio = (1332 - 440) / 700 = 1.27

Analysis: If this ratio is lower than industry average ratio then it leads to inefficiency in financial performance than industry avarage. As this ratio should be higher for better performance.

ii) Average collection period = 365 days / Accounts receivables turnover ratio

Here,

Accounts recievables turnover ratio = Net credit sales / Average accounts recievables

Accounts recievables turnover ratio = 3000 / ((560 + 600)/2) = 3000 / 580 = 5.17

Now, put the figures into formula,

Average collection period = 365/5.17 = 70.60 days

Analysis : If average collection period is lower than industry average then this is good & positive sign for receivables recovery. As lower ratio is preferred generally.

iii) Fixed assets turnover = Net sales / Average fixed assets

Fixed assets turnover = 3000 / ((1300 + 1400)/2)

Fixed assets turnover = 3000 / 1350 = 2.22

Analysis : If this ratio is lower than industry average then this shows negative performance of the firm. As higher ratio indicates greater efficiency.

iv) Operating margin profit = Operating profit(EBIT)/Sales = 500/3000 = 0.1667 or 16.67%

Analysis : Higher ratio is indication of greater profit margin. If this ratio is lower than industry average then this indicates the lower profit margin than industry average.

v) Times interest earned ratio = EBIT / Interest

Times interest earned ratio = 500 / 100 = 5

Analysis : If this ratio is lower than industry average then there is lower efficiency. As higher ratio indicates higher efficiency.

vi) Degree of financial leverage = EBIT / (EBIT - Interest)

Degree of financial leverage=500/(500 - 100)=1.25

Analysis : If this ratio is lower than industry average then it is positive sign which shows lower risk. As lower ratio indiactes lower risk burden.

b) Return on equity (ROE) using Dupont technique (2018)

= Profit margin * Total assets turnover * Equity multiplier

ROE using Dupont = (Net income/Net sales) * (Net sales /Total assets) * (Total assets / Total equity)

ROE using DuPont = (332/3000) * (3000/2732) * (2732/1142(note))

ROE using DuPont = 0.1107 * 1.0981 * 2.3923

ROE using DuPont = 0.2908 or 29.08%

Note: Total equity = Common stock + Additional paid up capital + retained earnings

Total equity = 100 + 200 + 842 = 1142